OpenStudy (luigi0210):

Graph the conic:

ganeshie8 (ganeshie8):

guess I would start by isolating \(y\) and find the horizontal/vertical asymptotes

OpenStudy (luigi0210):

We're suppose to rotate the coordinate the planes, I'm not sure how to do that..

OpenStudy (luigi0210):

*rotate the coordinate plane

OpenStudy (luigi0210):

So far I know it's a hyperbola by using the discriminant

ganeshie8 (ganeshie8):

yeah angle of rotation for eliminating xy term : \[\large \cot(2\theta) = \dfrac{A-C}{B}\] http://math.sci.ccny.cuny.edu/document/show/2685

ganeshie8 (ganeshie8):

\[\large \cot (2\theta) = \dfrac{0-0}{1} = 0 \implies \theta = \dfrac{\pi}{4}\]

ganeshie8 (ganeshie8):

\(\large x = x'\cos \theta - y' \sin \theta\) \(\large y = x'\sin \theta + y' \cos\theta\)